Suppose you are going to construct a circular fish pond in your garden. You want the pond to cover an area of 300ft^2. What is the radius of the pond?
Answer by nerdybill(7384)

Suppose you are going to construct a circular fish pond in your garden. You want the pond to cover an area of 300ft^2. What is the radius of the pond?
.
Since area of a circle is
(1/2)(pi)r^2
.
300 = (1/2)(pi)r^2
300 = (1/2)(3.14)r^2
600 = (3.14)r^2
600/3.14 = r^2
191.0828 = r^2
3.72 feet = r